OpenStudy (anonymous):

PLEASE HELP!!! ANSWER GETS MEDAL Solve for x: -2(x + 3) = -2(x + 1) - 4

OpenStudy (anonymous):

YOU WANT ANSURS?

OpenStudy (anonymous):

No I just want help please

OpenStudy (anonymous):

@SWAG

OpenStudy (anonymous):

When you simplify both sides you get -2x-6=-2x-6 0=0 I think the answer is all real numbers

OpenStudy (plainntall):

To solve for x you want to get all the terms on one side of the equation and 0 on the other side. -2(x + 3) = -2(x + 1) - 4 -2x-6=-2x-1-4 -2x-6=-2x-5 0=1 there is no solution for x where the equation is true. Are you sure the original equation is correct?

OpenStudy (plainntall):

I see, I made an error, it is true for all real numbers.

OpenStudy (anonymous):

What about this: @plainntall Solve for x: 4 over 5 x + 4 over 3 = 2x

OpenStudy (anonymous):

@plainntall

OpenStudy (anonymous):

If I understand what you've written 4/(5x+4) divided by 3=2x is this right

OpenStudy (anonymous):

No its like 4 4 - x + - = 2x 5 3

OpenStudy (anonymous):

@DSS